# Reinforcement Learning Discovers Efficient Decentralized Graph Path Search Strategies

### v1

@@ 1 Introduction
In this paper, we will study the problem of decentralized path graph search using local information. We will consider social networks and we will discuss how the proposed method can be directly applied to any networks for which topological and node attribute information is available. Indeed, prior experiments in human social networks, such as Stanley Milgram’s renowned “small world" experiment [12] [^1] The goal of Milgram’s experiment was to investigate the degree of connectedness among people in the United States, leading to the concept of “six degrees of separation”, the idea that any two people on Earth (i.e., the nodes in the global social network) are connected by a chain of no more than six acquaintances, which has also entered popular culture [13]. Milgram selected participants from Nebraska and Kansas. Each participant was given a letter and instructed to send it to a target person, a stockbroker living in Boston. However, they could only send the letter to someone they knew personally, characterized by certain (node) attributes who they thought might be closer to the target. Each recipient of the letter would then forward it to someone they knew personally, continuing this process until the letter reached the target. On average, it took about six steps for the letters to reach the stockbroker in Boston [14]. reveals the existence of short paths in social networks that are discoverable solely through local graph topology and high-level node attributes, e.g., characteristics of the individuals, such as their occupation, their high-school, and so on. Many social networks exhibit two key properties that make decentralized search with partial information possible and efficient. The first, homophily [15], reflects the tendency of individuals to connect with others who share similar attributes. The second is the heterogeneity of local structure in many networks, in which nodes are often organized into highly connected communities [16], with a smaller number of weak ties [17, 18] or central connectors [19] bridging these node clusters and acting as shortcuts.
@@ 2.2 Reinforcement Learning for Graph Routing and Search
One of the first instances of applying RL to such scenarios is the adaptive Q-routing algorithm [34]. The authors formulated the MDP with states as a tuple of the current node and destination node. They used a tabular Q-learning algorithm [35] where agents at each node on a graph maintain a lookup table of routing time estimates for neighbor - destination pairs. This method allows for adaptation in networks with changing topology using dynamic updates and does not require a central routing controller. Several extensions and variations on this idea have been explored [36, 37], but all suffer from the main pitfall of tabular RL methods: poor scalability.
Interest in learning to route has been reignited recently by several works that employ function approximation for better scalability. Notably, Graph Neural Networks (GNNs) [38] have emerged as a suitable learning representation for allowing RL policies to generalize to similar yet unseen conditions and graph topologies. In this line of work, Valadarsky et al. [39] considered learning routing policies on a dataset of historical traces of pairwise demands and applying them in new traffic conditions. The MDP is framed as learning a set of edge weights from which the routing strategy is determined. More recent work by Almasan et al. [40] leveraged a GNN representation trained using a policy gradient algorithm. They frame actions as the choice of a middle-point for a given flow given start and target nodes, with previous action choices becoming part of the state.
Another important line of work studies how to perform search on graphs. In contrast to routing, for search tasks there is no notion of a link load associated with traversing a particular node or edge in the graph. Notable contributions in this direction include work by Pándy et al. [21], where RL agents are tasked with learning a heuristic function for augmentation of A* search.
The problem of knowledge graph completion may also be viewed as graph traversal in instances with heterogeneous edge types [41] and with a target node that is not specified a priori. Das et al. [22] proposed an MDP formulation of this task, in which an agent chooses the next relationship to traverse given the current node. A proportion of the true relationships in the knowledge graph is masked and used to provide the reward signal for training the agent via REINFORCE. The M-Walk method [23] builds further in this direction by leveraging the determinism of the transition dynamics. Therefore, training with trajectories from a Monte Carlo Tree Search (MCTS) policy [42] can overcome the reward sparsity associated with the random exploration of model-free methods.
@@ 3.1.1 State & Action Space
We frame the search problem as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [43] taking place over an attributed, undirected, and unweighted graph structure $G=(V,E)$ with $n$ nodes and $m$ edges. An agent is placed on each node $u_{i}\in V$ in the graph, while the edges $E$ indicate direct bidirectional communication links between agents. An attribute vector $\mathbf{x}_{u_{i}}\in\mathbb{R}^{d}$ is associated with each agent. The aim is to find a path starting from an initial node $u_{\text{src}}$ to a designated target node $u_{\text{tgt}}$.
At each timestep $t$ of an MDP episode, a node $u_{i}$ receives a message $\mathbf{m}\in\mathbb{R}^{d}$ specifying the attributes $\mathbf{x}_{u_{\text{tgt}}}$ of the target node (but not its identity). It chooses as an action one of its neighbours, denoted $\mathcal{N}(u_{i})$, to pass the message on to. All other agents take no action at this step. We denote the presence or absence of a message at a given node $u_{i}$ at time $t$ with the binary indicator $M_{t}^{(i)}$.
To ensure MDP stationarity, the state of the receiver node at each step is conditioned on the current episode’s target node with variable $u_{\text{tgt}}$. We emphasize that this is a technical detail required to define the MDP, and the agent does not have visibility of the location of the target node beyond the attributes given in the message. The states and actions are therefore defined as:
$S_{t}=\{S_{t}^{(i)}|u_{i}\in V\}$ (1)
$S_{t}^{(i)}=(M_{t}^{(i)},u_{\text{tgt}})$ (2)
$A_{t}=\{A_{t}^{(i)}|u_{i}\in V\}$ (3)
$A_{t}^{(i)}|(M_{t}^{(i)}=1)=\mathcal{N}(u_{i})$ (4)
$A_{t}^{(i)}|(M_{t}^{(i)}=0)=\emptyset.$ (5)
3.1.2 Observation Space
@@
In accordance with our motivations, we provide agents with only local observations of the graph topology. Concretely, we equip each agent $u_{i}$ with observations of 1-hop ego subgraphs $G_{u_{i}}$ centered on its neighboring nodes, including visibility of pairwise edges between 1-hop neighbors. Symmetrically, the transmitted message comprises an ego graph centered on the target node. More formally, the observation provided to agent $i$ at time $t$ is defined as:
$O_{t}^{(i)}|(M_{t}^{(i)}=1)$ =($=(\{G_{u_{i}}|u\in\mathcal{N}(u_{i})\},G_{u_{\text{tgt}}})$ (6)
$O_{t}^{(i)}|(M_{t}^{(i)}=0)$ =$=\{G_{u_{i}}|u\in\mathcal{N}(u_{i})\},$ (7)
where
$G_{u_{i}}$ $=(V_{u_{i}},E_{u_{i}})$ (8)
$V_{u_{i}}$ $=\mathcal{N}(u_{i})\cup\{u_{i}\}$ (9)
$E_{u_{i}}$ =$=\{(u_{k},u_{l})\in E|u_{k}\in V_{u_{i}}\land u_{l}\in V_{u_{i}}\}.$ (10)
As it is possible to note from the definition above, the observation will be different depending on the absence or presence of the message on the node itself.
3.1.3 Transitions and Rewards
@@
If an action is permitted, the message moves deterministically to the selected node, updating states accordingly:
$M_{t+1}^{(i)}=$$\begin{cases}1\text{ if }M_{t}^{(j)}=1,A_{t}^{(j)}=u_{i},\\ 0\text{ otherwise.}\end{cases}$ (11)
@@
The episode ends when the message reaches the target node, yielding a collective reward of +1 for the agents and terminating the episode. Additionally, to prevent agents from becoming stuck in action cycles, we introduce a truncation criterion: after $T_{\text{max}}$ interaction steps with the environment, the episode ends without rewarding the agents. More formally:
$R_{t+1}^{(i)}=$$\begin{cases}1\text{ if }M_{t}^{(i)}=1,A_{t}^{(i)}=u_{\text{tgt}},\\ 0\text{ otherwise.}\end{cases}$ (12)
3.2 Model Description
@@
As it is common in the CDTE paradigm, we utilize parameter-sharing across agent networks. In the training scheme, a centralized agent receives localized observations from individual agents at each step, and is tasked with selecting optimal actions in the search path. The optimal decision is first learnt, and then replicated and distributed to individual nodes at execution time.
@@ 3.2.1 Policy Design
We formulate the choice of neighbor to which the message should be transmitted based on values output by an MLP-parameterized policy network $f_{\pi}$. The policy network is applied for each neighbor $u_{j}\in\mathcal{N}(u_{i})$ of the node $u_{i}$ that is currently in possession of the message at time $t$, and the SoftMax function is used to derive a probability distribution. Concretely, the policy is defined as:
$\pi(u_{j}|u_{i})=\frac{\text{exp}(f_{\pi}([\mathbf{x}_{u_{j}}||\mathbf{x}_{u_{\text{tgt}}}]))}{\sum_{u_{k}\in\mathcal{N}(u_{i})}{\text{exp}(f_{\pi}([\mathbf{x}_{u_{k}}||\mathbf{x}_{u_{\text{tgt}}}]))}},$ (13)
@@
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes concatenation. Similarly, to estimate the value function, we use the current node $u_{i}$ and the target node $u_{\text{tgt}}$ attributes that are passed through an MLP-parameterized value network $f_{v}$:
$v(u_{i},u_{\text{tgt}})=f_{v}([\mathbf{x}_{u_{i}}||\mathbf{x}_{u_{\text{tgt}}}]).$ (14)
@@
It is interesting to note that the node features that are used as input to the policy and value networks will impact the effectiveness of the learned policies. The simplest choice is to use the raw node features $\mathbf{x}_{u_{i}}$, and we denote the resulting algorithm as MLPA2C. We also consider the simplest extension to this model that minimally incorporates local graph topology by augmenting node attributes with node degrees, i.e., $\mathbf{x}_{u_{i}}^{\text{WD}}=[\mathbf{x}_{u_{i}}||\text{deg}(u_{i})]$. We refer to this as MLPA2CWD.
3.2.2 GARDEN
@@
Recall the “hidden metric" hypothesis discussed in Section 2.1, which posits that a viable policy can be motivated by moving through the graph to reduce node distance, provided a good approximation of the underlying metric is obtained. Instead of prescribing that the raw node attributes should be used to approximate this metric, we propose that relevant node features, which capture the potentially complex interplay between attributes and topologies, can be learned. To do so, we suggest replacing raw node attributes with learned embeddings $\mathbf{x}_{u_{i}}^{\text{GAT}}$ obtained from a Graph Attention Network (GAT) [25]. These embeddings are computed via a message-passing procedure starting from the raw node attributes and the graph topology.
@@
The method, which we refer to as Graph Attention-guided Routing for Decentralized Networks (GARDEN), is shown using pseudocode in Algorithm 1. The parameters of $f_{\text{rep}}$ are trained implicitly as we take gradient descent steps over the combined episodic loss $\sum_{t}L_{t}^{(\pi)}+L_{t}^{(v)}$. The node embeddings are recalculated at the start of each episode. The notation [[]]delimited-[]delimited-[][[$\cdot$]][ [ ⋅ ] ] denotes the partial stopping of gradients, and $H(p)$ denotes the entropy of a discrete distribution, given by $\sum_{i}p_{i}\text{log}(p_{i})$.
@@
1:  Input: Policy Network $f_{\pi}$, Value Network $f_{v}$, Graph Representation Network $f_{\text{rep}}$, Ego Graphs $\{G_{u}|u\in V\}$, entropy regularization coefficient $\lambda$, discount factor $\gamma$.
2:  Output: Learned policy, value and representation networks $f_{\pi},f_{v},f_{\text{rep}}$.
3:  for $i=1$ to $N_{\text{episodes}}$ do
4:     Initialize episode buffer $\mathcal{B}$
5:     Sample starting node $u$, target node $u_{\text{tgt}}$
6:     $\mathbf{m}=\mathbf{x}_{u_{\text{tgt}}}$
@@
8:        Compute $\mathbf{x}_{w}^{\text{GAT}}$ using $f_{\text{rep}}$
@@
10:     $t=0$
11:     while  $u\neq u_{\text{tgt}}$ and $t<T_{\text{max}}$ do
12:        Sample action $u^{\prime}\sim\pi(u;\mathbf{m})$
13:        Move message to node $u^{\prime}$, observe reward $r$
14:        Store transition $(u,u^{\prime},r)$ in $\mathcal{B}$
15:        $u\leftarrow u^{\prime}$
16:        $t=t+1$
17:     end while
18:     Initialize episode loss $L=0$
19:     for $(u,u^{\prime},r)$ in $\mathcal{B}$ do
20:        $\hat{A}=r+[[f_{v}([\mathbf{x}_{u^{\prime}}^{\text{GAT}}||\mathbf{m}])]]-f_{v}([\mathbf{x}_{u}^{\text{GAT}}||\mathbf{m}])$
21:        $L^{(\pi)}=-\hat{A}\log{f_{\pi}([\mathbf{x}_{u^{\prime}}^{\text{GAT}}||\mathbf{m}])}-\lambda H(\pi(u;\mathbf{m}))$
22:        $L^{(v)}=\hat{A}^{2}$
23:        $L=L+L^{(\pi)}+L^{(v)}$
24:     end for
25:     Take gradient descent step on L w.r.t. parameters of $f_{v},f_{\pi},f_{\text{rep}}$
26:  end for

### v2

@@ 1 Introduction
In this paper, we will study the problem of decentralized path graph search using local information. We will consider social networks and we will discuss how the proposed method can be directly applied to any networks for which topological and node attribute information is available. Indeed, prior experiments in human social networks, such as Stanley Milgram’s renowned “small world" experiment [12] [^1] Milgram’s experiment investigated the degree of connectedness among people in the United States, leading to the concept of “six degrees of separation”, the idea that any two people on Earth are connected by a chain of no more than six acquaintances, which has also entered popular culture [13]. Milgram selected participants from Nebraska and Kansas. Each participant was given a letter and instructed to send it to a target person, a stockbroker living in Boston. However, they could only send the letter to someone they knew personally, characterized by certain (node) attributes who they thought might be closer to the target. Each recipient of the letter would then forward it to someone they knew personally, continuing this process until the letter reached the target. On average, it took about six steps for the letters to reach the stockbroker in Boston [14]. reveals the existence of short paths in social networks that are discoverable solely through local graph topology and high-level node attributes, e.g., characteristics of the individuals, such as their occupation, their high-school, and so on. Many social networks exhibit two key properties that make decentralized search with partial information possible and efficient. The first, homophily [15], reflects the tendency of individuals to connect with others who share similar attributes. The second is the heterogeneity of local structure in many networks, in which nodes are often organized into highly connected communities [16], with a smaller number of weak ties [17, 18] or central connectors [19] bridging these node clusters and acting as shortcuts.
@@ 2.2 Reinforcement Learning for Graph Routing and Search
Early work on RL for routing demonstrated the potential of the MDP formalism [34, 35, 36], but suffered from the main pitfall of tabular RL methods: poor scalability. Interest has been reignited recently by several works that employ function approximation for scaling to larger problems. In this line of work, Valadarsky et al. [37] considered learning routing policies on a dataset of historical traces of pairwise demands and applying them in new traffic conditions. The MDP is framed as learning a set of edge weights from which the routing strategy is determined. Hope and Yoneki [38] expanded on this work by introducing a Graph Neural Network (GNN) [39] technique for function approximation, showing its advantages over using simple feedforward neural networks. More recent work by Almasan et al. [40] leveraged a GNN representation trained using a policy gradient algorithm. They frame actions as the choice of a middle-point for a flow given start and target nodes, with previous action choices becoming part of the state. Other recent works on RL for routing considered optimizing a weighted combination of delay and throughput [41] and deciding how to re-route the most important flows (i.e., those with the most traffic) given an initial routing scheme [42].
Another important line of work studies how to perform search on graphs. In contrast to routing, for search tasks there is no notion of a link load associated with traversing a particular node or edge in the graph. A notable contribution in this direction is work by Pándy et al. [21], where RL agents are tasked with learning a heuristic function for augmentation of A* search. Patankar et al. [43] considered the task of validating the way in which humans perform graph navigation, adopting two theories relating topological graph properties to minimizing gaps in knowledge or compressing existing knowledge. Their DQN-based agent parameterized by a GNN was validated successfully using human graph navigation trajectories.
The problem of knowledge graph completion may also be viewed as graph traversal in instances with heterogeneous edge types [44] and with a target node that is not specified a priori. Das et al. [22] proposed an MDP formulation of this task, in which an agent chooses the next relationship to traverse given the current node. A proportion of the true relationships in the knowledge graph is masked and used to provide the reward signal for training the agent via REINFORCE. The M-Walk method [23] builds further in this direction by leveraging the determinism of the transition dynamics. Therefore, training with trajectories from a Monte Carlo Tree Search (MCTS) policy [45] can overcome the reward sparsity associated with the random exploration of model-free methods. Zhang et al. [46] proposed a hierarchical method that features a high-level agent for choosing a cluster in which the target may be located, and a lower-level agent that navigates within the cluster.
@@ 3.1 MDP Formulation
We frame the search problem as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [47] taking place over an attributed, undirected, and unweighted graph structure $G=(V,E)$ with $n$ nodes and $m$ edges. An agent is placed on each node $u_{i}\in V$ in the graph, while the edges $E$ indicate direct bidirectional communication links between agents. An attribute vector $\mathbf{x}_{u_{i}}\in\mathbb{R}^{d}$ is associated with each agent. The aim is to find a path starting from an initial node $u_{\text{src}}$ to a designated target node $u_{\text{tgt}}$ by passing a single global message.
States. The global state $S_{t}$ at time $t$ is a tuple $\langle S_{t}^{(1)},S_{t}^{(2)},\dots,S_{t}^{(n)}\rangle$ composed of the states $S_{t}^{(i)}=(M_{t}^{(i)},u_{\text{tgt}})$ of the individual agents. Here, $M_{t}^{(i)}$ is an indicator variable that denotes the presence or absence of the message at a given node $u_{i}$ at time $t$. While specifying the target node $u_{\text{tgt}}$ is required for MDP stationarity, its identity is not provided to the agent in the observation.
[…]
@@
Observations. The environment emits a global observation $O_{t}=\langle O_{t}^{(1)},O_{t}^{(2)},\dots,O_{t}^{(n)}\rangle$ at each time step $t$, from which each node $u_{i}$ only observes its own component $O_{t}^{(i)}$. In accordance with our motivations, we provide agents with only local observations of the graph topology. Concretely, we equip each agent $u_{i}$ with observations of 1-hop ego subgraphs $G_{u_{j}}$ centered on its neighboring nodes, including visibility of pairwise edges between 1-hop neighbors. The observation will also contain information on the target node: if the agent possesses the message, it symmetrically can observe an ego graph centered on $u_{\text{tgt}}$. Formally, the observation $O_{t}^{(i)}$ is defined as $(\mathbf{m},\{G_{u_{j}}|u_{j}\in\mathcal{N}(u_{i}),G_{u_{\text{tgt}}}\})$ if $M_{t}^{(i)}=1$, and $\{G_{u_{j}}|u_{j}\in\mathcal{N}(u_{i})\}$ otherwise. The ego graph $G_{u_{j}}=(V_{u_{j}},E_{u_{j}})$ is defined such that $V_{u_{j}}=\mathcal{N}(u_{j})\cup\{u_{j}\}$ and $E_{u_{j}}=\{(u_{k},u_{l})\in E|u_{k}\in V_{u_{j}}\land u_{l}\in V_{u_{j}}\}$.
@@
Transitions. The message moves deterministically to the selected node, updating the indicator $M_{t+1}^{(i)}$ accordingly. Concretely, $M_{t+1}^{(i)}=1$ if $M_{t}^{(j)}=1\land A_{t}^{(j)}=a_{u_{j}\to u_{i}}$, and 00 otherwise.
@@
Rewards. The episode ends when the message reaches the target node, yielding a collective reward of +1 for the agents and terminating the episode. Formally, $R_{t+1}^{(k)}=1\ \forall k$ if $\exists i\ .\ M_{t}^{(i)}=1\land A_{t}^{(i)}=a_{u_{i}\to u_{\text{tgt}}}$, and 00 otherwise. To prevent agents from entering action cycles, we introduce a truncation criterion: the episode can also end after $T_{\text{max}}$ interaction steps with the environment.
3.2 Learning Architecture
@@
As it is common in the CTDE paradigm, we utilize parameter-sharing across agent networks. In the training scheme, a centralized agent receives localized observations from individual agents at each step, and is tasked with selecting optimal actions in the search path. The optimal decision is first learnt, and then replicated and distributed to individual nodes at execution time.
@@
The goal is to learn, for each agent, a policy $\pi_{u_{i}}$ that maps observations to actions. We formulate the choice of neighbor to which the message should be transmitted based on values output by an MLP-parameterized policy network $f_{\pi}^{\Theta_{1}}$. The policy network is applied for each neighbor $u_{j}\in\mathcal{N}(u_{i})$ of the node $u_{i}$ that is currently in possession of the message at time $t$, and the SoftMax function is used to derive a probability distribution. Concretely, for node $u_{i}$ in posession of the message $\mathbf{m}$, the policy is defined as:
$\pi_{u_{i}}(a_{u_{i}\to u_{j}}|O^{(i)})=\frac{\text{exp}(f_{\pi}^{\Theta_{1}}([\mathbf{x}_{u_{j}}||\mathbf{x}_{u_{\text{tgt}}}]))}{\sum_{u_{k}\in\mathcal{N}(u_{i})}{\text{exp}(f_{\pi}^{\Theta_{1}}([\mathbf{x}_{u_{k}}||\mathbf{x}_{u_{\text{tgt}}}]))}},$ (1)
@@
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes concatenation. Similarly, to estimate the value function, we pass the current node $u_{i}$ and the target node $u_{\text{tgt}}$ attributes through an MLP-parameterized value network $f_{v}^{\Theta_{2}}$:
$v(O^{(i)})=f_{v}^{\Theta_{2}}([\mathbf{x}_{u_{i}}||\mathbf{x}_{u_{\text{tgt}}}]),$ (2)
@@
Where $u_{i}$ is the node in possession of the message $\mathbf{m}$ at time $t$. It is interesting to note that the node features that are used as input to the policy and value networks will impact the effectiveness of the learned policies. The simplest choice is to use the raw node features $\mathbf{x}_{u_{i}}$, and we denote the resulting algorithm as MLPA2C. We also consider the simplest extension to this model that minimally incorporates local graph topology by augmenting node attributes with node degrees, i.e., $\mathbf{x}_{u_{i}}^{\text{WD}}=[\mathbf{x}_{u_{i}}||\text{deg}(u_{i})]$. We refer to this as MLPA2CWD.
3.3 GARDEN
@@
Recall the “hidden metric" hypothesis discussed in Section 2.1, which posits that a viable policy can be motivated by moving through the graph to reduce node distance, provided a good approximation of the underlying metric is obtained. Instead of prescribing that the raw node attributes should be used to approximate this metric, we propose that relevant node features, which capture the potentially complex interplay between attributes and topologies, can be learned. To do so, we suggest replacing raw node attributes with learned embeddings $\mathbf{x}_{u_{i}}^{\text{GAT}}$ obtained from a Graph Attention Network (GAT) [25], denoted $f_{\text{rep}}^{\Theta_{3}}$. These embeddings are computed via a message-passing procedure starting from the raw node attributes and the graph topology.
@@
The method, which we refer to as Graph Attention-guided Routing for Decentralized Networks (GARDEN), is shown using pseudocode in Algorithm 1. The full set of model parameters $\Theta=\{\Theta_{i}\}_{i=1}^{3}$ is trained implicitly as we take gradient descent steps over the combined episodic loss $\sum_{t}L_{t}^{(\pi)}+L_{t}^{(v)}$. The node embeddings are recalculated at the start of each episode. The notation [[]]delimited-[]delimited-[][[$\cdot$]][ [ ⋅ ] ] denotes the partial stopping of gradients, and $H(p)$ denotes the entropy of a discrete distribution, given by $\sum_{i}p_{i}\text{log}(p_{i})$.
@@
1:  Input: Policy Network $f_{\pi}^{\Theta_{1}}$, Value Network $f_{v}^{\Theta_{2}}$, Graph Representation Network $f_{\text{rep}}^{\Theta_{3}}$, Ego Graphs $\{G_{u}|u\in V\}$, entropy regularization coefficient $\lambda$, discount factor $\gamma$.
2:  Output: Learned policy, value and representation networks $f_{\pi}^{\Theta_{1}},f_{v}^{\Theta_{2}},f_{\text{rep}}^{\Theta_{3}}$.
3:  Randomly initialize model parameters $\Theta=\{\Theta_{i}\}_{i=1}^{3}$.
4:  for $i=1$ to $N_{\text{episodes}}$ do
5:     Initialize episode buffer $\mathcal{B}$
6:     Sample starting node $u$, target node $u_{\text{tgt}}$
@@
8:        Compute $\mathbf{x}_{w}^{\text{GAT}}$ using $f_{\text{rep}}^{\Theta_{3}}$
@@
10:     $\mathbf{m}=\mathbf{x}_{u_{\text{tgt}}}^{\text{GAT}}$
11:     $t=0$
12:     while  $u\neq u_{\text{tgt}}$ and $t<T_{\text{max}}$ do
13:        Sample action $a_{u\to u^{\prime}}\sim\pi_{u}(\cdot\text{ }|\mathbf{m})$
14:        Move message to node $u^{\prime}$, observe reward $r$
15:        Store transition $(u,u^{\prime},r)$ in $\mathcal{B}$
16:        $u\leftarrow u^{\prime}$
17:        $t=t+1$
18:     end while
19:     Initialize episode loss $L=0$
20:     for $(u,u^{\prime},r)$ in $\mathcal{B}$ do
21:        $\hat{A}=r+[[f_{v}^{\Theta_{2}}([\mathbf{x}_{u^{\prime}}^{\text{GAT}}||\mathbf{m}])]]-f_{v}^{\Theta_{2}}([\mathbf{x}_{u}^{\text{GAT}}||\mathbf{m}])$
22:        $L^{(\pi)}=-\hat{A}\log{f_{\pi}^{\Theta_{1}}([\mathbf{x}_{u^{\prime}}^{\text{GAT}}||\mathbf{m}])}-\lambda H(\pi(u|\mathbf{m}))$
23:        $L^{(v)}=\hat{A}^{2}$
24:        $L=L+L^{(\pi)}+L^{(v)}$
25:     end for
26:     Take gradient descent step on L w.r.t. $\Theta$
27:  end for